# Layer by Layer: Uncovering Hidden Representations in Language Models

### v1

@@ 4.4 Impact of Chain-of-Thought Finetuning
Figure 5: Token-level prompt entropy across sequence length for Qwen 2.5 and Qwen 2.5-Math models. The base model (Qwen 2.5) exhibits greater prompt compression, while the finetuned model (Qwen 2.5-Math) maintains higher entropy, indicating more information retention.
Recent work has highlighted Chain-of-Thought (CoT) finetuning as a powerful strategy for improving reasoning capabilities (Arefin et al., 2024; DeepSeek-AI, 2025). To examine its effects, in Figure 5 we compare Qwen 2.5 and Qwen 2.5-Math (Yang et al., 2024a, b), where the latter underwent additional math pretraining and CoT finetuning. Measuring token-level prompt entropy across sequence length reveals that the finetuned model maintains higher entropy with lower variance across examples.
@@ 5 Extreme Input Conditions
To gain a deeper understanding of the underline factors which effect the representation quality, we check how each layer responds to different types of inputs. We use Pythia 410M to three types of extreme prompts and measure prompt entropy across layers (Figure 6). We find that:
@@
    Token repetition compresses intermediate layers. As $p$ increases (i.e., more repeated tokens), prompt entropy decreases sharply in mid-depth layers. This indicates that the model effectively recognizes and encodes these repetitive patterns, discarding redundancy in its internal representation.
@@
    Random tokens inflate early-layer entropy. When we introduce token-level randomness, entropy increases significantly in the first few layers, revealing that these initial layers are especially sensitive to noise. By contrast, deeper layers appear more robust to such perturbations.
  3. 3.
    Prompt length raises raw entropy but grows sublinearly once normalized. Longer inputs naturally boost the unnormalized entropy because more tokens create more variation. However, normalized entropy expands at a slower rate, suggesting each additional token contributes less unique information.
Overall, these results confirm that intermediate layers play a major role in handling complex or unusual inputs, selectively compressing or filtering out repetitive patterns while retaining crucial distinctions. At the same time, early layers respond more sensitively to noise, and the incremental benefit of adding more tokens diminishes with prompt length. This behavior highlights the diverse ways in which different layers balance the trade-off between preserving and discarding information, further underscoring the unique strengths of intermediate representations.
@@ 6 Comparison to Vision Transformers
Although our focus has mainly been on language models, similar questions arise in computer vision. Vision architectures and training regimes differ widely, ranging from fully supervised methods to self-supervised approaches, and from bidirectional encoders to autoregressive transformers.
To investigate whether our findings generalize to vision models, we examine five representative vision approaches: ViT (Dosovitskiy et al., 2021), a supervised Transformer trained on labeled data; BEiT (Bao et al., 2022), a self-supervised encoder that reconstructs masked patches, analogous to masked token prediction in language; DINOv2 (Oquab et al., 2024), a self-supervised approach leveraging augmentations and exponential moving average teachers; MAE (He et al., 2022), a self-supervised framework that masks patches and reconstructs them, akin to masked autoencoders in language; and AIM (El-Nouby et al., 2024), an autoregressive Transformer that predicts the next patch in an image sequence (GPT-style next-token prediction). We evaluate each model on ImageNet-1k via layer-wise probing and our framework’s metrics.
A departure from language models for most vision Transformers.
Figure 14 shows that ViT, BEiT, DINOv2, and MAE exhibit strictly increasing downstream accuracy toward final layers, unlike language models. These models also show steadily increasing invariance metrics with depth, suggesting that without an autoregressive objective, vision Transformers have less need for drastic transformations at mid-depth.
@@
In contrast, AIM—which is explicitly autoregressive over image patches—shows an entropy “valley” and corresponding peak in downstream accuracy at its intermediate layers (El-Nouby et al., 2024). This mimics the patterns we observe in LLMs like Pythia, suggesting that autoregressive training induces an information bottleneck mid-depth. As in language modeling, forcing a strictly left-to-right (or patch-to-patch) prediction can drive the model to compress non-local details earlier, then re-expand relevant features.
@@
Taken together, these results indicate that the strong mid-layer compression observed in LLMs is not purely a property of “sequential token data” vs. “image patch data,” but rather a byproduct of autoregressive training. While various self-supervised (or fully supervised) objectives in vision often foster more uniform feature building across layers, autoregressive vision models develop the same mid-layer bottlenecks and sweet spots that we see in language. Thus, the architectural and objective design—especially whether or not a model is autoregressive—appears crucial in shaping layer-wise representation quality, regardless of domain.
@@ 7 Discussion and Conclusion
In this work, we investigate the representation quality of intermediate layers in LLMs, shedding light on their critical role in downstream task performance. We introduce a unified framework of evaluation metrics, establish theoretical connections among them, and apply these metrics to analyze Transformer-based architectures, SSMs, and vision models. One key phenomenon unveiled by prompt entropy was an information bottleneck in the middle layers of autoregressive transformers in both vision and language domains. Furthermore, our results reveal that intermediate layers often surpass final layers in representation quality, emphasizing their importance for feature extraction. DiME, curvature, and infoNCE correlate very well with downstream performance, suggesting a fundamental connection between representation and generalizability.
In conclusion, our study deepens the understanding of internal representation dynamics in LLMs. These insights not only enrich the theoretical foundations of model representations but also offer practical implications for optimizing model design, training strategies, and real-world applications. Future research could investigate the underlying causes of intermediate layer compression and develop specialized metrics tailored to LLMs, enabling more precise and effective representation evaluation.
@@
8 Acknowledgements
Oscar Skean is supported by the Office of the Under Secretary of Defense for Research and Engineering under award number FA9550-21-1-0227.
@@ Appendix B Discussion on Prompt Entropy
We follow the work of (Wei et al., 2024) and use $\alpha$-order matrix-based entropy (Giraldo et al., 2014; Skean et al., 2023, 2024), which serves as a tractable surrogate for traditional Rényi’s $\alpha$-order entropy (Rényi, 1961). The quantity is calculated using a similarity kernel $\kappa$ on a batch of samples drawn from a distribution, without making explicit assumptions on what the true distribution is. The choice of kernel $\kappa$ is flexible and can be any infinitely divisible kernel such as the Gaussian kernel, linear kernel, or Laplacian kernel, among others. For this work, we restrict ourselves to the linear kernel $\kappa(a,b)=ab^{T}$. This choice is motivated by the linear representation hypothesis (Park et al., 2024b) which finds that large language model representations encode high-level concepts such as truth (Burns et al., 2022), honesty (Mallen & Belrose, 2024), and part-of-speech (Mamou et al., 2020) in linearly separable manifolds.
@@ Appendix D Prompt Augmentations
Appendix E Extreme Prompts
E.1 Increasing Repetition
@@ Appendix F Theorems
(Random Unit Vectors are Nearly Orthogonal) Suppose we have $m$ unit vectors in $\mathbb{R}^{d}$, that are distributed according to the uniform distribution on the hyper-sphere. Then with probability at least $1-m^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}$, we have that for any pair $i,j$, $i\not=j$,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\epsilon.$
@@
Notice that the probability of not landing in the $\epsilon$ band $T_{\epsilon}\subset\mathbb{S}_{n-1}$ around the equator of the hypersphere of dimension $n$ can be bounded as,
$\mathbb{P}(T_{\epsilon}^{c})\leq\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}.$
Which is a result from (Wainwright, 2019). Notice that for sufficiently small epsilon, this means that the dot product of any two, randomly chosen vectors can be bounded with high probability. Indeed notice that,
$\arccos(\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle)=\theta,$
and that if $\mathbf{v_{j}}\in T_{\epsilon}$, then, treating $\mathbf{v_{i}}$ as $e_{1}$, the basis vector, without loss of generality, we have that,
$\theta\leq\arccos(\frac{\epsilon}{2}).$
So then,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\cos(\arccos(\frac{\epsilon}{2}))=\frac{\epsilon}{2}\leq\epsilon.$
@@
$\mathbb{P}(\exists i,j:\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle>\epsilon)$ $\leq\sum_{i\not=j}\mathbb{P}(\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle>\epsilon)$
$\leq m^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}.$
So then with probability at least $1-m^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2}}$, we have that, for any pair $i,j$,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\epsilon.$
@@
(Maximum Prompt Entropy implies Large Dataset Entropy.) Suppose we have a orthogonally equivarient representation model $Z$ such that for all sequences $Z_{i}=Z(X_{i})$ the prompt entropy is maximal and the rows are unit. Suppose also that the data distribution $\mathbf{Data}$ is a isotropic unit Gaussian. Suppose we draw sequences of length $L=D$ from the data distribution. Then with probability $1-N^{2}\sqrt{2\pi}e^{\frac{-n\epsilon^{2}}{2N^{2}}}$ over draw of $\{\mathbf{x_{i}}\}_{i=1}^{N}\sim\mathbf{Data}$, we have that,
$|e^{-S_{2}(QQ^{\top})}-\frac{N}{D^{2}}|\leq\epsilon$
@@
First note that, since the prompt entropy is maximal for each sample $Z(X_{i})$, then the matrix $K_{Z}=ZZ^{\top}$ is full rank. Since by assumption each row of $Z$ has unit rows, then we know that $\|Z\|_{F}^{2}=L=\sum_{k=1}^{L}\sigma_{k}^{2}$. In particular we also know that $\sigma_{i}=\sigma_{j}$ for all pairs $i,j$ by the assumption that the prompt entropy is maximized. In particular we then know that $ZZ^{\top}$ is a orthogonal matrix, and the rows of $Z$ form an orthonormal set. We can then write, for some $O_{i}$ a rotation matrix, that,
$\mathbf{q_{i}}=\frac{1}{D}\sum_{i=1}^{D}\mathbf{z_{i}}=\frac{1}{D}O_{i}\mathbf{1}.$
Since by assumption our model $Z(\cdot)$ is orthogonally equivarient, and the Data distribution is radially symmetric, it follows that these $\{\mathbf{q_{i}}\}_{i=1}^{N}$ are random points on the hypersphere of radius $\frac{1}{\sqrt{D}}$. This means that the matrix $\sqrt{D}Q$ consists of rows that are uniform points on hypersphere of radius $1$. Now notice that,
$\|QQ^{\top}\|_{F}^{2}$ $=\frac{1}{D^{2}}\|DQQ^{\top}\|_{F}^{2}$
=$\frac{1}{D^{2}}(\sum_{i=1}^{N}\|\sqrt{D}q_{i}\|^{2}+\sum_{i\not=j}\langle\sqrt{D}q_{i},\sqrt{D}q_{j}\rangle).$
Since $\sqrt{D}q_{i}$ is a unit vector this will simplify to,
$\|QQ^{\top}\|_{F}^{2}=\frac{1}{D^{2}}(N+\sum_{i\not=j}\langle\sqrt{D}q_{i},\sqrt{D}q_{j}\rangle).$
@@
$S_{2}(QQ^{\top})=\log\left(\frac{1}{\|QQ^{\top}\|_{F}^{2}}\right),$
we have that, $e^{-S_{2}(QQ^{\top})}=\|QQ^{\top}\|_{F}^{2}$. In particular,
$|e^{-S_{2}(QQ^{\top})}-\frac{N}{D^{2}}|\leq\epsilon.$
@@
(Minimal Prompt Entropy Implies Small Dataset Entropy 2) Suppose we have a orthogonally equivarient representation model $Z$ such that for all sequences $Z_{i}=Z(X_{i})$ the prompt entropy is minimal and the rows are unit. Suppose also that the data distribution $\mathbf{Data}$ is a isotropic unit Gaussian. Suppose we draw sequences from the data distribution. Then with probability $1-N^{2}\sqrt{2\pi}e^{\frac{-D^{3}\epsilon^{2}}{2N^{6}}}$ over the draw of $\{\mathbf{x_{i}}\}_{i=1}^{N}\sim\mathbf{Data}$, we have that,
$|e^{-S_{2}(QQ^{\top})}-\frac{m^{3}}{N^{2}}|\leq\epsilon.$
@@
Since the prompt entropy is minimal for each sample, we know that each $Z(X_{i})$ will be a rank one matrix, so we can write it as the outer product. In particular, we can write $Z(X_{i})=\mathbf{v_{i}}{\mathbf{u_{i}}}^{\top}$. However, since the rows of $Z(X_{i})$ are of unit length, we know that all the rows are identical, so we may write $Z(X_{i})=\mathbf{v_{i}}\mathbf{1}^{\top}$. Then, it follows that,
$\mathbf{q_{i}}=\frac{1}{D}\sum_{i=j}^{D}\mathbf{z_{j}^{i}}=\frac{N}{D}\mathbf{v}_{i}.$
In particular the matrix $\frac{D}{N}Q$ has rows that are all unit vectors, and these are randomly distributed uniformly on the hypersphere. Now notice that,
$\|QQ^{\top}\|_{F}^{2}$ =$\sum_{i=1}^{N}\|q_{i}\|^{2}+\sum_{i\not=j}\langle q_{i},q_{j}\rangle$
=$\sum_{i=1}^{N}\frac{N^{2}}{D^{2}}\|v_{i}\|^{2}+\sum_{i\not=j}\frac{N^{2}}{D^{2}}\langle v_{i},v_{j}\rangle$
=$\frac{N^{3}}{D^{2}}+\sum_{i\not=j}\frac{N^{2}}{D^{2}}\langle v_{i},v_{j}\rangle.$
Now by proposition, with probability at least $1-N^{2}\sqrt{2\pi}e^{\frac{-D^{3}\epsilon^{2}}{2N^{6}}}$, we know that, for all $i\not=j$,
$\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle\leq\frac{\epsilon D^{2}}{N^{3}}.$
So then,
$|\|QQ^{\top}\|_{F}^{2}-\frac{N^{3}}{D^{2}}|\leq\epsilon.$
@@
Appendix G Additional Plots & Visualizations

### v2

@@ 4.4 Impact of Chain-of-Thought Finetuning
Figure 5: Token-level prompt entropy across sequence lengths for Qwen 2.5 and Qwen 2.5-Math. The base model (Qwen 2.5) exhibits greater prompt compression, while the finetuned (Qwen 2.5-Math) has higher entropy, indicating more information retention.
Recent work has highlighted Chain-of-Thought (CoT) finetuning as a powerful strategy for improving reasoning capabilities (Arefin et al., 2025; DeepSeek-AI, 2025). To examine its effects on representations, in Figure 5 we compare Qwen 2.5 and Qwen 2.5-Math (Yang et al., 2024), where the latter underwent additional math pretraining and CoT finetuning. Measuring token-level prompt entropy across sequence length reveals that the finetuned model maintains higher entropy with lower variance across examples.
@@ 5 Extreme Input Conditions
To better probe the underlying factors affecting representation quality, we inspect each layer’s responsiveness to different input types. We use Pythia-410M on three types of extreme prompts and measure prompt entropy across layers (Figure 6). We prove examples of these prompts in Appendix F. Overall, we find that:
@@
    Token repetition compresses intermediate layers. As $p$ increases (i.e., more repeated tokens), prompt entropy decreases sharply in mid-depth layers, suggesting that the model recognizes/encodes repetitive patterns and discards redundancy in its internal representation.
@@
    Random tokens inflate early-layer entropy. Adding token-level randomness, increases entropy significantly in early layers, revealing their sensitivity to noise. In contrast, deeper layers are more robust.
Overall, these results confirm that intermediate layers play a major role in handling complex or unusual inputs, selectively compressing or filtering out repetitive patterns while retaining crucial distinctions. Early layers are more sensitive to noise and the incremental benefit of adding more tokens diminishes with prompt length. This behavior highlights the diverse ways in which different layers balance the trade-off between preserving and discarding information, underscoring the significance of intermediate representations.
@@ 6 Comparison to Vision Transformers
Do our findings extend to other domains like computer vision? Vision models employ diverse architectures and training objectives from fully supervised learning to self-supervised methods, and from bidirectional to autoregressive encoders. Their diversity provides an ideal testbed to examine how well our findings generalize and how different training objectives shape internal representations.
We examine several representative vision approaches: ViT (Dosovitskiy et al., 2021), a supervised transformer trained on labeled data; CLIP (Radford et al., 2021), a weakly supervised image encoder; BEiT (Bao et al., 2022), a self-supervised encoder that reconstructs masked patches; DINOv2 (Oquab et al., 2024), a self-supervised approach leveraging augmentations and exponential moving average teachers; MAE (He et al., 2022), a self-supervised approach that reconstructs images from masked patches; AIM (El-Nouby et al., 2024), an autoregressive transformer that predicts the next patch in an image sequence (GPT-style next-token prediction); and AIMv2 (Fini et al., 2025), which extends AIM with a multimodal next-token prediction task. In Figure 14, we evaluate every model layer on ImageNet-1k with attention probing and our suite of metrics.
@@
AIM, which predicts image patches sequentially, exhibits the same entropy "valley" and accuracy peak at intermediate layers that we observed in language models like Pythia. This pattern suggests that autoregressive training, whether over text tokens or image patches, consistently creates a mid-depth information bottleneck. The sequential prediction constraint forces models to compress non-local contextual information early in processing, then selectively re-expand the most relevant features for accurate prediction. AIM’s strong intermediate performance was first noted in (El-Nouby et al., 2024). Interestingly, while the AIMv2 model does not show improved intermediate accuracy, it still produces an entropy valley. We hypothesize this difference is due to the multimodal text-vision pretext task, which may alter information compression dynamics.
Vision transformers behave differently from language models.
All models except for AIM exhibit strictly increasing downstream accuracy toward final layers. Similar trends have been shown for ResNets (Sorscher et al., 2022), where few-shot classification error is strictly decreasing across layers. Most non-autoregressive vision models show steadily increasing dataset entropy. The notable exception is BEIT, which exhibits a substantial intermediate dip. Taken together, the results suggest that without an autoregressive objective, vision transformers have less need for drastic transformations at mid-depth.
@@
The strong mid-layer compression observed in LLMs seems to be not purely a property of “sequential token data” vs. “image patch data,” but rather a byproduct of pretraining. While various self-supervised (or fully supervised) objectives in vision foster more uniform feature building across layers, autoregressive vision models develop similar mid-layer bottlenecks that we see in language. Thus, the objective design–whether or not a model is autoregressive—appears crucial in shaping layer-wise representation quality, regardless of domain.
@@ 7 Discussion and Conclusion
We investigated the representation quality of intermediate layers in LLMs and their role in downstream task performance. We introduced a unified framework of evaluation metrics, establish theoretical connections among them, and apply these metrics to analyze transformer-based architectures, SSMs, and vision models. A key phenomenon unveiled by prompt entropy was an information bottleneck in the middle layers of autoregressive transformers in both vision and language domains. Furthermore, we show that intermediate layers often surpass final layers in representation quality, holding implications for feature relevance and extraction. DiME, curvature, and infoNCE correlate well with downstream performance, suggesting a fundamental connection between representation and generalizability.
In conclusion, our work studies the internal representation dynamics in LLMs, offering theoretical and empirical insights as well as practical implications for optimizing model design and training strategies. Future work should further investigate the underlying causes of intermediate layer compression and do explicit finetuning to control compression.
@@
Acknowledgements
We thank the anonymous reviewers for their valuable feedback, which helped improve the clarity and presentation of our results. We are also grateful to Diego Doimo, Artemii Novoselov, Jhoan Keider Hoyos Osorio, Luis Sanchez, and Matteo Saponati (listed alphabetically) for fruitful discussions and helpful pointers to related literature. Oscar Skean is supported by the Office of the Under Secretary of Defense for Research and Engineering under award number FA9550-21-1-0227.
@@ Appendix B Discussion on Prompt Entropy
We follow the work of (Wei et al., 2024) and use $\alpha$-order matrix-based entropy (Giraldo et al., 2014; Skean et al., 2023, 2024), which serves as a tractable surrogate for traditional Rényi’s $\alpha$-order entropy (Rényi, 1961). The quantity is calculated using a similarity kernel $\kappa$ on a batch of samples drawn from a distribution, without making explicit assumptions on what the true distribution is. The choice of kernel $\kappa$ is flexible and can be any infinitely divisible kernel such as the Gaussian kernel, linear kernel, or Laplacian kernel, among others. For this work, we restrict ourselves to the linear kernel $\kappa(a,b)=ab^{T}$. This choice is motivated by the linear representation hypothesis (Park et al., 2024b) which finds that large language model representations encode high-level concepts such as truth (Burns et al., 2023), honesty (Mallen & Belrose, 2024), and part-of-speech (Mamou et al., 2020) in linearly separable manifolds.
@@ Appendix D Prompt Augmentations
Appendix E Using Evaluation Metrics as a Performance Proxy
Model Supervised (Best) Unsupervised
Naive (Last) min-DiME min-infoNCE min-Dataset Entropy
Pythia-410M 52.0 45.5 48.5 46.2 48.1
LLM2Vec-8B 66.3 63.9 60.0 64.3 50.4
Table 2: Average MTEB Performance (%) across Different Layer Selection Schemes
We previously demonstrated strong correlations between our unsupervised evaluation metrics and downstream performance. These correlations can be exploited to select high-performing layers for a given task entirely without supervision, as suggested by prior work (Agrawal et al., 2022; Garrido et al., 2023; Thilak et al., 2024).
In Figure 2, we apply this unsupervised layer selection approach to Pythia-410M and LLM2Vec-8B using the 32-task MTEB benchmark introduced in Section 4.1. Rather than computing task accuracies for every layer, we compute DiME, infoNCE, and dataset entropy for each task across all layers in a single forward pass. For each task, we then select the layer that minimizes one of these metrics—leveraging their negative correlation with downstream performance.
This straightforward yet effective method yields substantial performance improvements with no supervision. For example, DiME-based layer selection boosts the average MTEB score of Pythia-410M by 3%.
Appendix F Extreme Prompts
F.1 Increasing Repetition
@@ Appendix G Theorems
(Random Unit Vectors are Nearly Orthogonal) Suppose we have $m$ unit vectors in $\mathbb{R}^{D}$, that are distributed according to the uniform distribution on the hyper-sphere. Then with probability at least $1-m^{2}\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2}}$, we have that for any pair $i,j$, $i\not=j$,
$|\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle|\leq\epsilon.$
@@
We can begin by defining the central $\epsilon$-band around a slice of the hypersphere $\mathbb{S}_{D-1}$ as,
$T_{\epsilon}=\{z\in\mathbb{S}_{D-1}:|\langle z,e_{1}\rangle|\leq\epsilon/2\},$
where $e_{1}$ denotes the first basis vector. The probability of a uniformly distributed vector on the unit sphere not landing in $T_{\epsilon}\subset\mathbb{S}_{D-1}$ can be bounded as,
$\mathbb{P}(T_{\epsilon})\geq 1-\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2}}.$
Now, treating $\mathbf{v_{i}}$ as $e_{1}$, the basis vector, without loss of generality, we have that, when $\mathbf{v_{i}},\mathbf{v_{j}}$ are uniformly distributed on the hyper-sphere,
$\mathbb{P}(|\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle|<\epsilon)\leq\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2}}$
@@
$\mathbb{P}(\exists i,j:\langle|\mathbf{v_{i}},\mathbf{v_{j}}\rangle|>\epsilon)$ $\leq\sum_{i\not=j}\mathbb{P}(|\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle|>\epsilon)$
$\leq m^{2}\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2}}.$
So then with probability at least $1-m^{2}\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2}}$, we have that, for any pair $i,j$,
$|\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle|\leq\epsilon.$
@@
(Maximum Prompt Entropy implies Large Dataset Entropy.) Suppose we have a orthogonally equivarient representation model $Z$ such that for all sequences $Z_{i}=Z(X_{i})$ the prompt entropy is maximal and the rows are unit. Suppose also that the data distribution $\mathbf{Data}$ is a isotropic unit Gaussian. Suppose we draw sequences of length $L=D$ from the data distribution. Then with probability $1-N^{2}\sqrt{2\pi}e^{\frac{-D\epsilon^{2}}{2N^{2}}}$ over draw of $\{\mathbf{x_{i}}\}_{i=1}^{N}\sim\mathbf{Data}$, we have that,
$|e^{-S_{2}(QQ^{\top})}-\frac{N}{L^{2}}|\leq\epsilon$
@@
First note that, since the prompt entropy is maximal for each sample $i$, which we denote $Z_{i}=Z(X_{i})$, then the matrix $K_{Z}=Z_{i}Z_{i}^{\top}$ is full rank. Since by assumption each row of $Z_{i}$ has unit rows, then we know that $\|Z_{i}\|_{F}^{2}=L=\sum_{k=1}^{L}\sigma_{k}^{2}$. In particular we also know that $\sigma_{l}=\sigma_{j}$ for all pairs $l,j$ by the assumption that the prompt entropy is maximized. In particular we then know that $Z_{i}Z_{i}^{\top}$ is a orthogonal matrix, and the rows of $Z_{i}$ form an orthonormal set. We can then write, for some $O_{i}$ a rotation matrix, that,
$\mathbf{q_{i}}=\frac{1}{L}\sum_{i=1}^{L}\mathbf{z_{i}}=\frac{1}{L}O_{i}\mathbf{1}.$
We will denote the average over sequences of length $L$, across all $N$ samples, by the dataset matrix $\bar{Z}=(\mathbf{q}_{1},\mathbf{q}_{2},\ldots\mathbf{q}_{N})^{\top}$. Since by assumption our model $Z(\cdot)$ is orthogonally equivariant, and the Data distribution is radially symmetric, it follows that these $\{\mathbf{q_{i}}\}_{i=1}^{N}$ are random points on the hypersphere of radius $\frac{1}{\sqrt{L}}$. This means that the matrix $\sqrt{D}\bar{Z}$ consists of rows that are uniform points on hypersphere of radius $1$. Now notice that,
$\|\bar{Z}\bar{Z}^{\top}\|_{F}^{2}$ $=\frac{1}{L^{2}}\|L\bar{Z}\bar{Z}^{\top}\|_{F}^{2}$
=$\frac{1}{L^{2}}(\sum_{i=1}^{N}\|\sqrt{L}q_{i}\|^{2}+\sum_{i\not=j}\langle\sqrt{L}q_{i},\sqrt{L}q_{j}\rangle).$
Since $\sqrt{L}q_{i}$ is a unit vector this will simplify to,
$\|\bar{Z}\bar{Z}^{\top}\|_{F}^{2}=\frac{1}{L^{2}}(N+\sum_{i\not=j}\langle\sqrt{L}q_{i},\sqrt{L}q_{j}\rangle).$
@@
$S_{2}(\bar{Z}\bar{Z}^{\top})=\log\left(\frac{1}{\|\bar{Z}\bar{Z}^{\top}\|_{F}^{2}}\right),$
we have that, $e^{-S_{2}(\bar{Z}\bar{Z}^{\top})}=\|\bar{Z}\bar{Z}^{\top}\|_{F}^{2}$. In particular,
$|e^{-S_{2}(\bar{Z}\bar{Z}^{\top})}-\frac{N}{L^{2}}|\leq\epsilon.$
@@
(Minimal Prompt Entropy Implies Small Dataset Entropy.) Suppose we have a orthogonally equivariant representation model $Z$ such that for all sequences $Z_{i}=Z(X_{i})$ the prompt entropy is minimal and the rows are unit. Suppose also that the data distribution $\mathbf{Data}$ is a isotropic unit Gaussian. Suppose we draw sequences from the data distribution. Then with probability $1-N^{2}\sqrt{2\pi}e^{\frac{-D^{3}\epsilon^{2}}{2N^{8}}}$ over the draw of $\{\mathbf{x_{i}}\}_{i=1}^{N}\sim\mathbf{Data}$, we have that,
$|e^{-S_{2}(\bar{Z}\bar{Z}^{\top})}-\frac{N^{3}}{L^{2}}|\leq\epsilon.$
@@
Since the prompt entropy is minimal for each sample, we know that each $Z(X_{i})$ will be a rank one matrix, so we can write it as the outer product. In particular, we can write $Z(X_{i})=\mathbf{v_{i}}{\mathbf{u_{i}}}^{\top}$. However, since the rows of $Z(X_{i})$ are of unit length, we know that all the rows are identical, so we may write without loss of generality, $Z(X_{i})=\mathbf{v_{i}}\mathbf{1}^{\top}$. Then, it follows that,
$\mathbf{q_{i}}=\frac{1}{L}\sum_{i=j}^{L}\mathbf{z_{j}^{i}}=\frac{N}{L}\mathbf{v}_{i}.$
We will write the dataset average matrix as before as $\bar{Z}=(\mathbf{q}_{1},\mathbf{q}_{2},\ldots\mathbf{q}_{N})^{\top}$. In particular the matrix $\frac{D}{N}\bar{Z}$ has rows that are all unit vectors, and these are randomly distributed uniformly on the hyper-sphere. Now notice that,
$\|\bar{Z}\bar{Z}^{\top}\|_{F}^{2}$ =$\sum_{i=1}^{N}\|\mathbf{q}_{i}\|^{2}+\sum_{i\not=j}\langle\mathbf{q}_{i},\mathbf{q}_{j}\rangle$
=$\sum_{i=1}^{N}\frac{N^{2}}{L^{2}}\|\mathbf{v}_{i}\|^{2}+\sum_{i\not=j}\frac{N^{2}}{L^{2}}\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$
=$\frac{N^{3}}{L^{2}}+\sum_{i\not=j}\frac{N^{2}}{L^{2}}\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle.$
Now by the prior proposition, with probability at least $1-N^{2}\sqrt{2\pi}e^{\frac{-D^{3}\epsilon^{2}}{2N^{8}}}$, we know that, for all $i\not=j$,
$|\langle\mathbf{v_{i}},\mathbf{v_{j}}\rangle|\leq\frac{\epsilon L^{2}}{N^{4}}.$
So then we have that,
$|\|\bar{Z}\bar{Z}^{\top}\|_{F}^{2}-\frac{N^{3}}{L^{2}}|\leq\sum_{i\not=j}\frac{N^{2}}{L^{2}}\langle|\mathbf{v}_{i},\mathbf{v}_{j}\rangle|\leq\frac{1}{L^{2}}\sum_{i\not=j}\epsilon\leq\epsilon.$
@@
Appendix H Additional Plots & Visualizations